40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Aug 29 01:52:22 2017, gautam, Update, SUS, Test mass actuator calibration ALS_comparison.pdf
    Reply  Tue Aug 29 15:04:59 2017, gautam, Update, SUS, ETMY Oplev PIT loop gain changed ETMY_OLPIT.pdfGTR_comparison.pdf
    Reply  Wed Oct 18 17:34:09 2017, gautam, Update, SUS, ASDC 
Message ID: 13392     Entry time: Wed Oct 18 17:34:09 2017     In reply to: 13265
Author: gautam 
Type: Update 
Category: SUS 
Subject: ASDC 

Summary:

The signal path for the ASDC signal is AS55 PD --> D990543 (interface board) --> D990694 (whitening board) --> D000076 (AA board) --> ADC Ch 31. Everything in this signal chain should be able to handle signals in the range +/- 10V, which should correspond to the full range of our +/-10V, 16bit ADCs. But the ASDC signal seems to saturate at ~2000 counts (i.e. turning up the analog whitening gain doesn't make the signal get any bigger than this). I investigated this a little more today.

Details:

  • The ASDC signal is derived from the AS55 photodiode. According to the schematic, the Op27 that supplies this voltage is powered by +/- 15V, so the output should be able to swing between at least +/- 12V.
  • The DC signal goes from the DB15 connector on the side of the PD to the LSC electronics rack, 1Y2, where it is interfaced with an LSC PD Interface Card, D990543. Again, per the schematic, the Op27 driving this voltage is powered by +/- 15V, and so the available output voltage swing should be greater than +/-12V.
  • The D990543 output is to its backplane connector. There is an adaptor board hooked up to the backplane that makes these outputs available to a LEMO connector. A LEMO-SMA cable then pipes this output to a D990694.
    • I decided to test the functionality of this board.
    • Disconnected the SMA ASDC input signal (CH8 on the board).
    • Drove that channel with an SR function generator and gradually turned up the Vpp of the input signal (sine wave at 145Hz).
    • Monitored the ASDC channel on dataviewer while doing this.
    • Saw that the ASDC signal saturated at ~2000 counts. Turning up the signal amplitude did not have any effect.
  • From the whitening board, the signal goes through an anti-aliasing module (D000076). The final stage LT1125s on these boards should also be supplied with +/-15V.

So the problem lies somewhere downstream of the D990694. There are other anomalous behaviours of this channel - e.g. engaging the analog whitening filters changes the DC offset of the signal. I am going to pull out this board to check it out.

Why does this matter? I want to calibrate the ASDC level (and eventually the other PD DC signals as well) into Watts. This is useful for IFO diagnostics, noise budgeting the shot noise level etc.

According to the AS55 schematic, the DC transimpedance is 66.7 ohms. I claim that the DC power on the AS55 photodiode during a DRMI (no arms) lock is ~1mW. The C30642 photodiode (InGaAs) responsivity is ~0.8 A/W. So I'd expect ~50mV to be the signal level into the ADC (assuming gain of all the other electronics in the signal chain at the start of this elog is unity). This corresponds to ~163 counts (since the ADC conversion factor is 2^16 counts over 20volts). The DC signal level I observed is ~200 counts. So things seem roughly consistent.

*Note: Despite my above statement, I don't think it is true that the AS110 PD has more light on it - the BS splitting the light between

AS55 and AS110 PDs is a 50-50 BS, and using the crude method of putting an Ophir power meter in front of both PDs and

monitoring the power while the Michelson was swinging around freely showed roughly the same maximum value.

ELOG V3.1.3-